x^2-9y^2+36y-45=0
1
Expert's answer
2019-07-25T13:39:47-0400

x2-9y2+36y-45=0, x2-(9y2-36y+36)-45+36=0, x2-(3y-6)2-9=0, x2-9(y-2)2=9, x2/9-(y-2)2=1. We have the equation of hyperbola here.